OpenStudy (samara8954):

Martha estimated there were 93 marbles in a jar for a contest. The actual number of marbles in the jar was 115. What was the percent error of Martha's estimation? 80.87% 23.66% 19.13% 22%

OpenStudy (danjs):

Take the difference in the values, divided by the total marbles in the jar

OpenStudy (danjs):

multiply that by 100

OpenStudy (danjs):

She was (115-93) off the value, out of 115 actual marbles.

OpenStudy (jhannybean):

Percent error = \(\dfrac{\left|\text{actual - theoretical}\right|}{\text{theoretical}} \times 100\%\)

OpenStudy (danjs):

\[\frac{ 115 - 93 }{ 115 } *100 = ~~~??~~percent\]